# HelixVS: Deep Learning Enhanced Structure-based Virtual Screening Platform for Hit Discovery

### v1

@@ Abstract
Drug discovery through virtual screening (VS) has become a popular strategy for identifying hits against protein targets. VS has the potential to reduce the cost and time needed for manual selection and wet-laboratory experiments, thereby accelerating drug discovery pipelines. Improving the cost-effectiveness of virtual screening is a significant challenge, aiming to explore larger compound libraries while maintaining lower screening costs. Here, we present HelixVS, a structure-based VS platform enhanced by deep learning models. HelixVS integrates a precise deep learning-based pose-scoring model and a pose-screening module into a multi-stage VS process, enabling more effective screening of active compounds. Compared to classic molecular docking tools like Vina, HelixVS demonstrated significantly improved screening performance across nearly a hundred targets, achieving an average 2.6-fold higher enrichment factor (EF) and more than 10 times faster screening speed. We applied HelixVS in four drug development pipelines, targeting both traditional competitive drug-binding pockets and novel protein-protein interaction interfaces. Wet-lab validations across these pipelines consistently identified active compounds, with over 10% of the molecules tested in wet labs demonstrating activity at M or even nM levels. This demonstrates the ability of HelixVS to identify high-affinity ligands for various targets and pockets. Furthermore, we provide a publicly available and free version of HelixVS with limited computing power to assist drug development scientists in accelerating their drug discovery processes. The HelixVS online service is available at: https://paddlehelix.baidu.com/app/drug/helixvs/forecast. We also support private deployment solutions to meet the data security requirements of pharmaceutical companies, research institutes, and other organizations. For collaboration inquiries, please contact: baidubio_cooperate@baidu.com.
@@ 1 Introduction
Based on this understanding, we developed HelixVS, a cost-effective virtual screening platform to enhance screening efficiency and success rate in early-stage drug discovery. HelixVS performs multi-stage screening, comprehensively taking advantage of both molecular docking tools and deep learning-based models to screen out potential active molecules against target proteins. HelixVS demonstrated significantly improved screening performance compared to using only molecular docking tools. Compared to Vina across over a hundred targets in the DUD-E dataset Mysinger et al. (2012), HelixVS can find an average of 159% more active molecules and run up to nearly 15 times faster. These results highlight the effectiveness of HelixVS’s multi-stage screening process.
@@
Superior Hit Rate: HelixVS significantly improves virtual screening hit rates through multi-stage screening strategies and integration of deep learning models. Testing on the DUD-E dataset shows that HelixVS can achieve enrichment factors (EF) of 44.205 and 26.968 at 0.1% and 1% respectively, significantly outperforming other methods. In multiple actual drug development projects, including difficult scenarios such as dual-target and protein-protein interface (PPI) binding, HelixVS has successfully identified active molecules. These results demonstrate HelixVS platform’s exceptional capability in identifying active compounds.
@@ 2.1 Multi-staged virtual screens with HelixVS
The HelixVS platform is a comprehensive drug discovery pipeline consisting of three stages to identify potential active molecules that can bind with a specific protein target (as shown in Figure 1).
@@ 4 Conclusion
Our studies have demonstrated that HelixVS not only enhances the efficiency of the screening process but also significantly improves the success rate by leveraging both classical molecular docking tools and advanced deep learning-based models. This dual approach ensures a comprehensive exploration of chemical space and a higher likelihood of discovering viable hit compounds. Furthermore, the platform’s ability to handle vast libraries and deliver rapid screening results has been instrumental in accelerating the early stages of drug development.

### v2

# HelixVS: Deep Learning–Enhanced Structure-Based Platform for Screening and Design

@@ Abstract
Drug discovery through virtual screening (VS) has become a popular strategy for identifying hits against protein targets. Alongside VS, molecular design further expands accessible chemical space. Together, these approaches have the potential to reduce the cost and time needed for manual selection and wet-laboratory experiments, thereby accelerating drug discovery pipelines. Improving the cost-effectiveness of virtual screening is a significant challenge, aiming to explore larger compound libraries while maintaining lower screening costs. Here, we present HelixVS, a structure-based VS platform enhanced by deep learning models. HelixVS integrates a precise deep learning-based pose-scoring model and a pose-screening module into a multi-stage VS process, enabling more effective screening of active compounds. Compared to classic molecular docking tools like Vina, HelixVS demonstrated significantly improved screening performance across nearly a hundred targets, achieving an average 2.6-fold higher enrichment factor (EF) and more than 10 times faster screening speed. We applied HelixVS in four drug development pipelines, targeting both traditional competitive drug-binding pockets and novel protein-protein interaction interfaces. Wet-lab validations across these pipelines consistently identified active compounds, with over 10% of the molecules tested in wet labs demonstrating activity at M or even nM levels. This demonstrates the ability of HelixVS to identify high-affinity ligands for various targets and pockets. In addition, the HelixVS platform has been extended with HelixVS-Syn, which enables design of novel compounds from reference scaffolds. These designed molecules are seamlessly integrated into the HelixVS screening workflow, allowing researchers to explore both existing chemical libraries and novel chemical space with high affinity, synthetic accessibility, and structural novelty. Furthermore, we provide a publicly available and free version of HelixVS with limited computing power to assist drug development scientists in accelerating their drug discovery processes. The HelixVS online service is available at: https://paddlehelix.baidu.com/app/drug/helixvs/forecast. We also support private deployment solutions to meet the data security requirements of pharmaceutical companies, research institutes, and other organizations. For collaboration inquiries, please contact: baidubio_cooperate@baidu.com.
@@ 1 Introduction
Based on this understanding, we developed HelixVS, a cost-effective virtual screening platform to enhance screening efficiency and success rate in early-stage drug discovery. HelixVS performs multi-stage screening, comprehensively taking advantage of both molecular docking tools and deep learning-based models to screen out potential active molecules against target proteins. Moreover, the platform is extended with HelixVS-Syn, a module that enables scaffold-based molecular design, further broadening the accessible chemical space before screening. HelixVS demonstrated significantly improved screening performance compared to using only molecular docking tools. Compared to Vina across over a hundred targets in the DUD-E dataset Mysinger et al. (2012), HelixVS can find an average of 159% more active molecules and run up to nearly 15 times faster. These results highlight the effectiveness of HelixVS’s multi-stage screening process.
@@
Superior Hit Rate: HelixVS significantly improves virtual screening hit rates through multi-stage screening strategies and integration of deep learning models. Testing on the DUD-E dataset shows that HelixVS can achieve enrichment factors (EF) of 44.205 and 26.968 at 0.1% and 1% respectively, significantly outperforming other methods. In multiple actual drug development projects, including difficult scenarios such as dual-target and protein-protein interface (PPI) binding, HelixVS has successfully identified active molecules.HelixVS-Syn further complements this capability. Under the same computational budget, HelixVS-Syn designs molecules that surpass those from the VS in synthetic accessibility, binding affinity, and novelty. These results demonstrate HelixVS platform’s exceptional capability in identifying active compounds.
@@ 2 The HelixVS Platform
The HelixVS platform supports multiple sources of molecular inputs, including its built-in large-scale compound library, user-defined molecular libraries, and newly designed molecules generated by HelixVS-Syn from reference scaffolds. All these inputs can be seamlessly integrated into the subsequent three-stage drug discovery pipeline, ensuring that diverse molecular sources undergo the same systematic screening to identify potential active compounds against specific protein targets.(as shown in Figure 1).
@@ 2.3 Compound Library Pre-processing
2.4 Design Novel Molecules with HelixVS-Syn
HelixVS-Syn enables design of novel compounds from a known molecular scaffold and integrates these designs into a complete virtual-screening workflow. It is a pathway-driven molecular design algorithm that starts from synthesizable and purchasable compounds. By integrating retrosynthetic knowledge with synthesis-guided modifications, it generates new molecules that remain synthetically accessible. This enables exploration of a broad chemical space while reducing computational cost and research time compared with large-scale screening.
We conducted a comparative experiment against HelixVS performed on a compound library of 820,000 molecules, using CDK5 as the target protein. From the 50,000 molecules generated by HelixVS-Syn, we selected the top 50 for statistical analysis.
Compared to HelixVS, HelixVS-Syn produces compounds with higher synthetic accessibility, affinity and novelty(how these metrics are calculated is reported in Appendix Section1.3).
| Method | Synthetic accessibility | Affinity | Novelty |
| HelixVS-Syn | 0.641 | 60.69 | 0.708 |
| HelixVS | 0.615 | 64.54 | 0.658 |
@@ 4 Conclusion
Our studies have demonstrated that HelixVS not only enhances the efficiency of the screening process but also significantly improves the success rate by leveraging both classical molecular docking tools and advanced deep learning-based models. This dual approach ensures a comprehensive exploration of chemical space and a higher likelihood of discovering viable hit compounds. Furthermore, the platform’s ability to handle vast libraries and deliver rapid screening results has been instrumental in accelerating the early stages of drug development.Notably, the integration of HelixVS-Syn further extends this capability by enabling scaffold-based molecular design, thereby expanding the accessible chemical space with novel, synthetic accessibility, and active candidates.
@@ 1.2 The Bimolecular Fluorescence Complementation Assay
1.3 Molecule Synthetic accessibility, Affinity and Novelty Evaluation Metric
Synthetic accessibility was evaluated by the Retro*Chen et al. (2020) success rate. The Retro* success rate is defined as the proportion of molecules in a set for which Retro* can identify a synthetic route. Affinity was assessed using calculated affinity scores.
where : total number of molecules under evaluation, : indicator function: if Retro* identifies a valid synthetic route for molecule , and otherwise.
Affinity was predicted by the DTA models.
Novelty was assessed by computing the Tanimoto similarity between the generated molecules and a reference compound library of 200,000 molecules. For each generated molecule, the mean similarity score to its 50 most similar molecules in the library was calculated. The overall novelty score was then obtained as the mean of these similarity values across all generated molecules, and defined as
where : total number of generated molecules, : Tanimoto similarity between the -th molecule and its -th most similar compound in the reference library : mean similarity of the -th molecule to its 50 most similar library molecules, : mean similarity across all generated molecules.